OpenStudy (anonymous):

A man walking from Ash Street to Bay treat covered 1/12 of the distance in 10 minutes. 1.) in 7/12 of an hour, what fraction of the distance will he walk? 2.) How many how will it take him to cover the distance?

OpenStudy (mathmale):

Seems to me that figuring out the man's speed (walking speed) is the way to start out. Let d be the distance from Ash to Bay. How would you express, algebraically, the distance from Ash to Bay? Divide your result by "10 minutes." What you'll have now will be the man's speed. Simplify the fractions, if you wish.

OpenStudy (texaschic101):

(1/12) / 10 = (12/12)/x --- 1/12 in 10 min = 12/12 in x min cross multiply 1/12x = (10)(12/12) 1/12x = 120/12 x = (120/12) / (1/12) x = 120/12 * 12/1 x = 120 minutes or 2 hrs So it will take him 120 minutes (or 2 hrs) to cover the total distance. Is this right mathmale ?

OpenStudy (mathmale):

Looks good. So now, what's the walker's speed?

OpenStudy (texaschic101):

7/12 of an hour = 7/12 * 60 = 420/12 = 35 minutes (1/12) / 10 = x/35 --- 1/12 in 10 min = x in 35 min cross multiply 10x = (1/12)(35) 10x = 35/12 x = (35/12) / 10 x = 35/12 * 1/10 x = 35/120 = 7/24 He will walk 7/24 of the distance in 7/12 of an hour You better get mathmale here to check this....I am not even sure if it is correct....there might even be an easier way to do this.

OpenStudy (mathmale):

I've suggested that we find the man's walking speed, if d represents the distance that he walks. His speed is \[\frac{ distance }{ time }=\frac{ \frac{ d }{ 12 } }{ 10~\min }=\frac{ d }{ 120~\min }=\frac{ d }{ 2~hr }\]

OpenStudy (mathmale):

How far will the man walk in 7/12 of one hour? Just multiply his rate by this time:\[(rate)(time)=\frac{ d }{ 2~hr }\frac{ 7~hr }{ 12 }\]

OpenStudy (mathmale):

So this result is identical to that of texaschic101. Nice work! Unfortunately, neither of us did anything to get @MoeshaKiel involved, and thus have earned 10 demerits apiece. @MoeshaKiel : Please share with us your "take" on this discussion. What's clear for you and what is not?

OpenStudy (texaschic101):

the way you did it looks easier....proportions can sometimes be confusing

OpenStudy (mathmale):

@MoeshaKiel : We are not told what the "distance" is, but we can certainly represent it by " d ". We ARE told that this guy can walk one twelfth of this distancd d in 10 minutes. How would you find the man's walking speed from this information?